# DGS-Net: Distillation-Guided Gradient Surgery for CLIP Fine-Tuning in AI-Generated Image Detection

### v1

@@ Abstract
The rapid progress of generative models such as GANs and diffusion models has led to the widespread proliferation of AI-generated images, raising concerns about misinformation, privacy violations, and trust erosion in digital media. Although large-scale multimodal models like CLIP offer strong transferable representations for detecting synthetic content, fine-tuning them often induces catastrophic forgetting, which degrades pre-trained priors and limits cross-domain generalization. To address this issue, we propose the Distillation-guided Gradient Surgery Network (DGS-Net), a novel framework that preserves transferable pre-trained priors while suppressing task-irrelevant components. Specifically, we introduce a gradient-space decomposition that separates harmful and beneficial descent directions during optimization. By projecting task gradients onto the orthogonal complement of harmful directions and aligning with beneficial ones distilled from a frozen CLIP encoder, DGS-Net achieves unified optimization of prior preservation and irrelevant suppression. Extensive experiments on 50 generative models demonstrate that our method outperforms state-of-the-art approaches by an average margin of 6.6%, achieving superior detection performance and generalization across diverse generation techniques.
@@ 1 Introduction
The rapid advancement of deep learning has driven significant progress in generative models such as generative adversarial networks (GANs) [karras2018progressive, park2019semantic, huang2024gan] and diffusion models [ho2020denoising, rombach2022high, wu2024infinite]. Owing to their low generation cost and high visual fidelity, these models have been widely applied across various domains, including digital art, entertainment, and content creation. However, their widespread use has also raised serious concerns about privacy violations, misinformation dissemination, and the erosion of trust in digital media. As a result, developing reliable detectors to verify the authenticity of AI-generated images has become an urgent and critical task. Despite notable progress in recent detection approaches [zhao2021multi, wu2024local, tan2024frequency, li2025pay], achieving robust generalization to unseen generative techniques remains a fundamental and unsolved challenge.
@@
Large-scale multimodal pre-trained models, such as CLIP [radford2021learning], have recently emerged as powerful tools for detecting AI-generated images. By leveraging transferable representations learned from massive image–text datasets, CLIP constructs a semantic embedding space with strong open-set scalability and cross-domain generalization [tan2025c2p, yan2024orthogonal], providing a robust foundation for identifying synthetic images [ojha2023towards, koutlis2024leveraging, zhang2025towards] produced by diverse generative models. However, CLIP’s pre-training objective is not explicitly optimized to capture generation artifacts—key cues for distinguishing authentic and synthetic content. To address this limitation, recent studies have adapted CLIP for detection tasks through either full fine-tuning or parameter-efficient tuning strategies [liu2024forgery, tan2025c2p, fu2025exploring] such as LoRA [hu2022lora]. Despite these efforts, fine-tuning often induces catastrophic forgetting, eroding the transferable priors and geometric structure of the pre-trained embedding space. Consequently, the adapted model may overfit to dataset-specific patterns rather than learning artifact-level features that generalize across different generative mechanisms.
@@
To investigate this phenomenon, we construct four datasets using ProGAN [karras2018progressive], R3GAN [huang2024gan], SDXL [podell2023sdxl], and SimSwap [chen2020simswap], each containing both real and generated images across multiple categories. We then extract features using three variants of CLIP: the original pre-trained model [ojha2023towards], a LoRA fine-tuned model, and our proposed method. The extracted features are visualized using t-SNE, as shown in Figure 1. The frozen CLIP model preserves the geometric structure and local coherence learned during pre-training but exhibits weak real/fake separability. In contrast, LoRA fine-tuning enhances separability but collapses the representation geometry and degrades cross-domain detection performance. These observations indicate that catastrophic forgetting in existing fine-tuning strategies compromises the transferable priors of the pre-trained model.
@@
To address these issues, we introduce a knowledge distillation framework [hinton2015distilling] tailored for CLIP-based detectors. As illustrated in Figure 1, the pre-trained CLIP features inherently contain numerous components unrelated to generation artifacts. Conventional feature distillation methods, which enforce global alignment between teacher and student representations, may inadvertently retain these task-irrelevant factors. Therefore, our goal is to preserve transferable pre-trained priors while suppressing task-irrelevant components. Specifically, we introduce a gradient-space decomposition strategy, where the positive half-space of task gradients is defined as harmful directions and the negative half-space as beneficial directions (see Sec.˜3 for details). Building on this principle, we propose a Distillation-guided Gradient Surgery Network (DGS-Net). During backpropagation, we first project the image gradients of the training network onto the orthogonal complement of the harmful directions estimated from text gradients, effectively suppressing components irrelevant to the detection task. Meanwhile, we exploit the beneficial descent directions extracted from a frozen CLIP image encoder to guide lightweight alignment of the training network’s representations, thereby preserving the transferable priors acquired during pre-training. As shown in Figure 1, our approach achieves strong real/fake discrimination while maintaining the pre-trained model’s prior knowledge, ultimately enhancing generalization across diverse generative models.
@@ 2.1 AI-Generated Image Detection
With the rapid advancement of generative models such as GANs [karras2018progressive, karras2021alias, sauer2022stylegan, huang2024gan] and diffusion models [ho2020denoising, dhariwal2021diffusion, li2024photomaker, wu2024infinite], the visual quality of synthetic images has improved dramatically, making it increasingly difficult for the human eye to distinguish real images from generated ones. In response to the potential misuse of these generative techniques, extensive research has been devoted to AI-generated image detection, which can be broadly categorized into artifact-based methods [tan2023learning, zheng2024breaking, tao2025sagnet, yan2025ns] and data-centric methods [wang2020cnn, chen2024drct, chen2025dual]. Artifact-based approaches typically exploit low-level forensic cues that expose generation traces, such as frequency irregularities [luo2021generalizing, tan2024frequency, yan2025generalizable], pixel-level inconsistencies [chen2020manipulated, chai2020makes, cavia2024real], gradient patterns [tan2023learning], neighboring-pixel dependencies [tan2024rethinking], and reconstruction errors [wang2023dire, cazenavette2024fakeinversion, chu2025fire, chen2025dual]. For example, BiHPF [jeong2022bihpf] enhances artifact visibility by applying dual high-pass filters; LGrad [tan2023learning] leverages gradients from pretrained CNNs to reveal universal forgery traces; NPR [tan2024rethinking] reinterprets upsampling operations to construct an efficient artifact representation; and DIRE [wang2023dire] captures generation discrepancies via reconstruction error analysis. In contrast, data-centric methods aim to learn more generalizable artifact representations through large-scale data and augmentation strategies. Representative efforts include data augmentation [wang2020cnn, li2025improving] and dataset construction [yu2024semgir, chen2024drct, guillaro2025bias, chen2025dual]. CNN-Spot [wang2020cnn] enhances model robustness by combining diverse augmentation techniques with large-scale GAN-generated datasets. SAFE [li2025improving] integrates cropping and augmentations such as ColorJitter, RandomRotation, and RandomMask to improve generalization. DRCT [chen2024drct] leverages diffusion-based reconstruction to achieve better semantic alignment, while B-Free [guillaro2025bias] mitigates dataset bias using self-conditioned inpainted reconstructions and content-aware augmentation.
@@
With the rapid rise of large pre-trained models, particularly vision–language models such as CLIP [radford2021learning], recent studies have begun exploiting their powerful feature extraction and semantic understanding capabilities for AI-generated image detection. The impressive generalization ability of CLIP demonstrated in UnivFD [ojha2023towards] has inspired a series of follow-up works that adapt or fine-tune CLIP for improved detection performance. Existing approaches can be broadly divided into two categories: (1) feature-based methods, which freeze the pre-trained network and optimize only the extracted features to enhance generalization [ojha2023towards, koutlis2024leveraging, zhang2025towards]; and (2) fine-tuning-based methods, which update model parameters to learn task-specific features through full fine-tuning, low-rank adaptation (LoRA) [liu2024mixture, tan2025c2p, yan2025ns], or lightweight adaptation modules [liu2024forgery, yan2024orthogonal, tao2025sagnet]. For instance, $\text{D}^{3}$ [yang2025d] utilizes CLIP as a powerful feature backbone and introduces a parallel branch to disentangle universal fake artifacts from multi-generator images. VIB-Net [zhang2025towards] applies a Variational Information Bottleneck to decouple and filter CLIP-extracted features, effectively suppressing task-irrelevant information. C2P-CLIP [tan2025c2p] injects category-consistent cues into the text encoder and employs contrastive learning to improve cross-modal alignment. Effort [yan2024orthogonal] constructs two orthogonal subspaces to preserve pre-trained priors while learning forgery-related representations. NS-Net [yan2025ns] further constructs a NULL-Space of semantic features to remove semantic interference embedded in visual features.
@@ 2.2 Knowledge Distillation
Knowledge Distillation [hinton2015distilling] is a powerful technique for transferring knowledge from a large teacher model to a smaller student model. Numerous studies [li2017mimicking, wang2019distilling, dai2021general, jia2024mssd] focus on selecting the most valuable features for knowledge distillation in downstream tasks, with the goal of preserving the richest prior knowledge. This technique has been successfully applied across a variety of domains, including visual recognition [kim2018paraphrasing, heo2019comprehensive, li2021online, li2023curriculum, huang2024etag], language model compression [jiao2019tinybert, park2021distilling, wu2023ad], and multimodal representation learning [fang2021compressing, chen2023enhanced, chen2024comkd]. A seminal work [chen2017learning] introduced the first distillation framework for object detection, combining feature imitation with prediction mimicking. [pfeiffer2020adapterfusion] incorporated a distillation loss into adapter-based fine-tuning to transfer representations from pre-trained models, preventing feature drift during adaptation. [wei2023improving] proposed feature-level knowledge distillation during fine-tuning, which helps the student model maintain cross-modal semantic alignment while learning downstream tasks.
However, our research shows that not all pre-trained knowledge is beneficial for downstream adaptation, as some representations can negatively impact task performance. In response, we extend the traditional distillation paradigm to selectively retain useful priors while suppressing irrelevant, harmful representations. Unlike conventional feature distillation methods [tian2019contrastive, miles2021information, chen2021distilling, yang2021knowledge], which penalize discrepancies at the representation level, our approach operates in the gradient space. It suppresses gradients that are irrelevant to the task while injecting a small, beneficial knowledge prior into the descent direction, ultimately enhancing the performance of AI-generated image detection.
@@ 3 Preliminaries
Before introducing our method, we first establish a directional convention for interpreting the gradient dynamics of classification losses. For any differentiable classification loss, the first-order gradient of the representation indicates the direction of the steepest local variation in the loss landscape [kingma2014adam]. Specifically, a positive gradient value at a given coordinate implies that a positive perturbation along that direction increases the loss (i.e., a harmful direction), whereas a negative gradient value implies that a positive perturbation decreases the loss (i.e., a beneficial direction).
@@ 4.1 Notation and Setup
Given image–label pair $(x,y)$ with $y\in\{0,1\}$, where $y=1$ denotes a fake image and $y=0$ denotes a real image, CLIP provides an image encoder $E_{\mathrm{img}}$ and a text encoder $E_{\mathrm{text}}$. The corresponding feature extraction processes are defined as:
@@
where $\theta$ denotes the trainable parameters of the image encoder, and $\phi$ is frozen since the text encoder is only used to extract semantic representations. Here, $z$ represents the text description associated with image $x$. In this work, we incorporate LoRA (Low-Rank Adaptation) layers as the trainable parameters $\theta$ for fine-tuning CLIP’s image encoder, while using BLIP [li2022blip] to automatically generate the corresponding text description $z$. Additionally, we maintain a frozen image encoder $E_{\mathrm{img}}^{T}$—a copy of the CLIP image encoder prior to fine-tuning—to extract reference features:

### v2

@@ Abstract
The rapid progress of generative models such as GANs and diffusion models has led to the widespread proliferation of AI-generated images, raising concerns about misinformation and trust erosion in digital media. Although large-scale multimodal models like CLIP offer strong transferable representations for detecting synthetic content, fine-tuning them often induces catastrophic forgetting, which degrades pre-trained priors and limits cross-domain generalization. To address this issue, we propose the Distillation-guided Gradient Surgery Network (DGS-Net), a novel framework that preserves transferable pre-trained priors while suppressing task-irrelevant components. Specifically, we introduce a gradient-space decomposition that separates harmful and beneficial descent directions during optimization. By projecting task gradients onto the orthogonal complement of harmful directions and aligning with beneficial ones distilled from a frozen CLIP encoder, DGS-Net achieves unified optimization of prior preservation and irrelevant suppression. Extensive experiments on 50 generative models demonstrate that our method outperforms state-of-the-art approaches by an average margin of 6.6%, achieving superior detection performance and generalization across diverse generation techniques.
Machine Learning, ICML
@@ 1 Introduction
The rapid advancement of deep learning has driven significant progress in generative models such as generative adversarial networks (GANs) (karras2018progressive; park2019semantic; huang2024gan) and diffusion models (ho2020denoising; rombach2022high; wu2024infinite). Owing to their low generation cost and high visual fidelity,they have been widely applied across various domains, including digital art, entertainment, and content creation. However, their widespread use has also raised serious concerns about privacy violations, misinformation dissemination, and the erosion of trust in digital media. As a result, developing reliable detectors to verify the authenticity of AI-generated images has become an urgent task. Despite notable progress in recent detection methods (zhao2021multi; tan2024frequency; li2025pay), achieving robust generalization to unseen generative techniques remains an unsolved challenge.
@@
Large-scale multimodal pre-trained models, such as CLIP (radford2021learning), have recently emerged as powerful tools for detecting AI-generated images. By leveraging transferable representations learned from massive image–text datasets, CLIP constructs a semantic embedding space with strong open-set scalability and cross-domain generalization (tan2025c2p; yan2024orthogonal), providing a robust foundation for identifying synthetic images (ojha2023towards) produced by diverse generative models. However, CLIP’s pre-training objective is not explicitly optimized to capture generation artifacts. To address this limitation, recent studies have adapted CLIP for detection tasks through parameter-efficient tuning strategies (liu2024forgery; tan2025c2p; fu2025exploring) such as LoRA (hu2022lora). Despite these efforts, fine-tuning often induces catastrophic forgetting, eroding the transferable priors and geometric structure of the pre-trained embedding space. Consequently, the adapted model may overfit to dataset-specific patterns rather than learning artifact-level features that generalize across different generative mechanisms.
@@
To investigate this phenomenon, we construct four datasets using ProGAN (karras2018progressive), R3GAN (huang2024gan), SDXL (podell2023sdxl), and SimSwap (chen2020simswap), each containing both real and generated images across multiple categories. We then extract features using three variants of CLIP: the original model (ojha2023towards), a LoRA fine-tuned model, and our proposed method. The extracted features are visualized using t-SNE, as shown in Figure 1. The frozen CLIP model preserves the geometric structure and local coherence learned during pre-training but exhibits weak real/fake separability. In contrast, LoRA fine-tuning enhances separability but collapses the representation geometry and degrades cross-domain detection performance. These observations indicate that catastrophic forgetting in existing fine-tuning strategies compromises the transferable priors of the pre-trained model.
@@
To address these issues, we introduce a knowledge distillation framework (hinton2015distilling) tailored for CLIP-based detectors. As illustrated in Figure 1, the pre-trained CLIP features inherently contain numerous components unrelated to generation artifacts. Conventional feature distillation methods, which enforce global alignment between teacher and student representations, may inadvertently retain these task-irrelevant factors. Therefore, our goal is to preserve transferable pre-trained priors while suppressing task-irrelevant components. Specifically, we introduce a gradient-space decomposition strategy, where the positive half-space of task gradients is defined as harmful directions and the negative half-space as beneficial directions (see Section˜3 for details). Building on this principle, we propose a Distillation-guided Gradient Surgery Network (DGS-Net). During backpropagation, we first project the image gradients of the training network onto the orthogonal complement of the harmful directions estimated from text gradients, effectively suppressing components irrelevant to the detection task. Meanwhile, we exploit the beneficial descent directions extracted from a frozen CLIP image encoder to guide lightweight alignment of the training network’s representations, thereby preserving the transferable priors acquired during pre-training. As shown in Figure 1, our approach achieves strong real/fake discrimination while maintaining the pre-trained model’s prior knowledge, ultimately enhancing generalization across diverse generative models.
@@ 2.1 AI-Generated Image Detection
With the rapid advancement of generative models such as GANs (karras2018progressive; karras2021alias; huang2024gan) and diffusion models (ho2020denoising; dhariwal2021diffusion; wu2024infinite), making it increasingly difficult for the human eye to distinguish real images from generated ones. Extensive research has been devoted to AI-generated image detection, which can be broadly categorized into artifact-based methods (zheng2024breaking; tao2025sagnet; yan2025ns) and data-centric methods (wang2020cnn; chen2025dual). Artifact-based approaches typically exploit low-level forensic cues that expose generation traces, such as frequency irregularities (luo2021generalizing; tan2024frequency; yan2025generalizable), pixel-level inconsistencies (chen2020manipulated; cavia2024real), gradient patterns (tan2023learning), neighboring-pixel dependencies (tan2024rethinking), and reconstruction errors (wang2023dire; chu2025fire; chen2025dual). For example, BiHPF (jeong2022bihpf) enhances artifact visibility by applying dual high-pass filters; NPR (tan2024rethinking) reinterprets upsampling operations to construct an efficient artifact representation; while DIRE (wang2023dire) captures generation discrepancies via reconstruction error analysis. In contrast, data-centric methods aim to learn more generalizable artifact representations through large-scale data and augmentation strategies. Representative efforts include data augmentation (wang2020cnn; li2025improving) and dataset construction (chen2024drct; guillaro2025bias; chen2025dual). SAFE (li2025improving) integrates cropping and augmentations such as ColorJitter, RandomRotation, and RandomMask to improve generalization, while B-Free (guillaro2025bias) mitigates dataset bias using self-conditioned inpainted reconstructions and content-aware augmentation.
@@
With the rapid rise of large pre-trained models, particularly vision–language models such as CLIP (radford2021learning), recent studies have begun exploiting their powerful feature extraction and semantic understanding capabilities for AI-generated image detection. The impressive generalization ability of CLIP demonstrated in UnivFD (ojha2023towards) has inspired a series of follow-up works, which can be broadly divided into two categories: (1) feature-based methods, which freeze the pre-trained network and optimize only the extracted features to enhance generalization (ojha2023towards; koutlis2024leveraging; zhang2025towards); and (2) fine-tuning-based methods, which update model parameters to learn task-specific features through full fine-tuning, low-rank adaptation (LoRA) (liu2024mixture; tan2025c2p; yan2025ns), or lightweight adaptation modules (liu2024forgery; yan2024orthogonal; tao2025sagnet). For instance, VIB-Net (zhang2025towards) applies a Variational Information Bottleneck to decouple and filter CLIP-extracted features, effectively suppressing task-irrelevant information. C2P-CLIP (tan2025c2p) injects category-consistent cues into the text encoder and employs contrastive learning to improve cross-modal alignment. Effort (yan2024orthogonal) constructs two orthogonal subspaces to preserve pre-trained priors while learning forgery-related representations. NS-Net (yan2025ns) further constructs a NULL-Space of semantic features to remove semantic interference embedded in visual features.
@@ 2.2 Knowledge Distillation
Knowledge Distillation (hinton2015distilling) is a powerful technique for transferring knowledge from a large teacher model to a smaller student model. Numerous studies (li2017mimicking; wang2019distilling; dai2021general; jia2024mssd) focus on selecting the most valuable features for knowledge distillation in downstream tasks, with the goal of preserving the richest prior knowledge. This technique has been successfully applied across a variety of domains, including visual recognition (kim2018paraphrasing; heo2019comprehensive; li2021online; li2023curriculum; huang2024etag), language model compression (jiao2019tinybert; park2021distilling; wu2023ad), and multimodal representation learning (fang2021compressing; chen2023enhanced; chen2024comkd). A seminal work (chen2017learning) introduced the first distillation framework for object detection, combining feature imitation with prediction mimicking. (pfeiffer2020adapterfusion) incorporated a distillation loss into adapter-based fine-tuning to transfer representations from pre-trained models, preventing feature drift during adaptation. (wei2023improving) proposed feature-level knowledge distillation during fine-tuning, which helps the student model maintain cross-modal semantic alignment while learning downstream tasks.
However, our research shows that not all pre-trained knowledge is beneficial for downstream adaptation, as some representations can negatively impact task performance. In response, we extend the traditional distillation paradigm to selectively retain useful priors while suppressing irrelevant, harmful representations. Unlike conventional feature distillation methods (tian2019contrastive; miles2021information; chen2021distilling; yang2021knowledge), which penalize discrepancies at the representation level, our approach operates in the gradient space. It suppresses gradients that are irrelevant to the task while injecting a small, beneficial knowledge prior into the descent direction, ultimately enhancing the performance of AI-generated image detection.
@@ 3 Preliminaries
Before introducing our method, we first establish a directional convention for interpreting the gradient dynamics of classification losses. For any differentiable classification loss, the first-order gradient of the representation indicates the direction of the steepest local variation in the loss landscape (kingma2014adam). Specifically, a positive gradient value at a given coordinate implies that a positive perturbation along that direction increases the loss (i.e., a harmful direction), whereas a negative gradient value implies that a positive perturbation decreases the loss (i.e., a beneficial direction).
@@ 4.1 Notation and Setup
Given image–label pair $(x,y)$ with $y\in\{0,1\}$, where $y=1$ denotes a fake image and $y=0$ denotes a real image, CLIP provides an image encoder $E_{\mathrm{img}}$ and a text encoder $E_{\mathrm{text}}$. The corresponding feature extraction processes are defined:
@@
where $\theta$ denotes the trainable parameters of the image encoder, and $\phi$ is frozen since the text encoder is only used to extract semantic representations. Here, $z$ represents the text description associated with image $x$. In this work, we incorporate LoRA (Low-Rank Adaptation) layers as the trainable parameters $\theta$ for fine-tuning CLIP’s image encoder, while using BLIP (li2022blip) to automatically generate the corresponding text description $z$. Additionally, we maintain a frozen image encoder $E_{\mathrm{img}}^{T}$—a copy of the CLIP image encoder prior to fine-tuning—to extract reference features:
@@
Figure 3: Classification using only text descriptions achieves an accuracy of approximately 60%. We used BLIP to convert images into textual descriptions and trained the detector directly on these text inputs. The resulting detection accuracy fluctuated around 60%, indicating that semantic information is partially correlated with the labels. However, most of them act as distractors that hinder cross-generator generalization.
Table 1: Cross-model Accuracy (Acc.) Performance on the AIGCDetectBench (zhong2023patchcraft) Dataset. The first column represents the accuracy of detecting real images (R.Acc.), and the others are the accuracy of detecting fake images (F.Acc.).
Detection Method Real Image Generative Adversarial Networks Other Diffusion Models mAcc.
CNN-Spot (wang2020cnn) 99.0 95.3 18.7 1.8 36.5 22.0 2.5 23.1 1.1 29.3 55.9 55.6 1.8 4.8 5.2 27.6 0.7 4.5 29.0
UnivFD (ojha2023towards) 92.3 98.9 90.5 79.2 55.7 48.7 91.1 96.9 92.8 26.9 96.3 96.0 12.7 75.6 61.2 84.7 45.6 62.3 72.7
FreqNet (tan2024frequency) 89.9 99.4 57.1 51.0 75.1 67.5 9.9 88.4 95.4 35.8 99.9 99.8 37.7 78.9 80.8 98.0 34.1 88.8 71.7
NPR (tan2024rethinking) 99.3 98.9 29.3 16.5 67.1 58.7 1.7 6.5 7.9 0.1 100.0 99.9 26.5 69.2 71.0 97.7 15.4 89.8 53.1
Ladeda (cavia2024real) 99.8 99.7 7.2 29.0 95.6 98.3 4.9 0.0 19.2 0.1 100.0 99.9 27.3 79.7 88.8 97.9 15.5 92.4 58.6
AIDE (yan2024sanity) 93.0 95.3 88.0 96.9 89.6 97.0 90.0 97.0 42.9 9.2 99.8 99.7 92.4 98.7 63.7 98.9 90.1 98.9 85.6
C2P-CLIP* (tan2025c2p) 99.8 100.0 99.8 99.8 72.2 78.0 87.7 100.0 34.9 67.6 100.0 99.9 50.5 73.3 28.4 99.7 93.3 98.1 82.4
DFFreq (yan2025generalizable) 98.0 98.0 70.9 93.6 99.0 98.9 93.8 97.4 3.5 76.0 99.9 99.8 65.9 87.8 94.8 99.1 76.0 95.9 85.2
SAFE (li2025improving) 99.2 99.7 85.6 83.5 88.1 96.7 89.0 99.9 8.4 17.3 99.8 99.6 36.8 90.5 86.3 98.5 84.0 92.0 80.8
Effort (yan2024orthogonal) 99.8 100.0 100.0 100.0 79.6 82.1 100.0 100.0 99.9 76.6 100.0 99.9 53.1 81.2 44.2 99.9 96.3 75.8 88.1
NS-Net (yan2025ns) 97.7 100.0 99.9 100.0 95.7 98.6 98.5 100.0 68.9 60.6 100.0 99.9 85.3 97.2 77.4 100.0 98.9 98.8 93.2
Table 2: Cross-model Average Precision (A.P.) Performance on the AIGCDetectBench (zhong2023patchcraft) Dataset.
CNN-Spot(wang2020cnn) 99.9 91.0 59.2 94.8 94.2 86.1 82.5 65.5 74.2 97.6 97.6 66.6 83.1 80.4 88.9 57.0 87.8 84.3
UnivFD (ojha2023towards) 99.9 96.2 94.7 90.9 91.1 97.6 99.7 83.0 83.5 99.3 99.1 65.2 95.3 91.7 97.4 87.2 89.8 91.9
FreqNet (tan2024frequency) 100.0 97.5 87.9 95.9 94.2 54.0 100.0 57.0 88.5 99.9 99.9 73.9 94.1 94.4 99.1 74.9 92.3 88.4
NPR (tan2024rethinking) 100.0 98.6 79.4 98.6 99.6 71.3 97.6 65.5 92.5 100.0 99.9 74.1 97.4 97.8 99.9 81.5 99.3 91.4
Ladeda (cavia2024real) 100.0 99.0 89.5 100.0 100.0 96.4 90.3 93.6 92.6 95.5 95.8 84.8 96.6 93.5 93.0 84.5 93.9 94.0
AIDE (yan2024sanity) 99.6 98.2 91.6 99.4 99.9 74.1 99.7 90.8 66.5 99.9 99.9 99.4 99.9 90.3 99.9 98.9 99.9 94.5
C2P-CLIP* (tan2025c2p) 100.0 100.0 99.9 99.3 99.7 98.8 100.0 99.6 98.3 100.0 100.0 96.3 98.9 90.4 100.0 99.9 100.0 98.8
DFFreq (yan2025generalizable) 100.0 98.7 98.4 100.0 100.0 98.1 99.8 89.0 87.5 100.0 99.9 98.7 99.5 99.7 100.0 99.3 99.9 98.1
SAFE (li2025improving) 100.0 99.5 97.8 99.9 100.0 96.6 100.0 73.8 91.0 100.0 100.0 98.1 99.9 99.9 100.0 99.9 100.0 97.4
Effort (yan2024orthogonal) 100.0 100.0 100.0 99.3 99.2 100.0 100.0 91.1 97.9 100.0 100.0 97.5 99.8 98.9 100.0 100.0 99.9 99.0
NS-Net (yan2025ns) 100.0 100.0 99.7 99.8 99.9 97.0 100.0 99.5 95.2 100.0 100.0 100.0 99.8 97.0 100.0 100.0 100.0 99.2